OpenStudy (anonymous):

Solve x over 2 - 3 = 7. a,5 b,10 c,20 d,40

OpenStudy (anonymous):

x/2 - 3=7 add 3 on both sides and then multiply both sides by 2...

OpenStudy (anonymous):

x/2-3=7 x/-1=7 x=-7

OpenStudy (anonymous):

X/2 - 3 =7 +3 +3 X/2 = 10 *2 *2 X = 20